# Non-Linear Scoring Model for Translation Quality Evaluation

### v2

@@ Abstract
Analytic Translation Quality Evaluation (TQE), based on Multidimensional Quality Metrics (MQM), has traditionally used a linear error-to-penalty scale calibrated to a reference sample size of 1000100020002000 words. However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition.
Building on the previously introduced Multi-Range framework [GladkoffEtAl2024], this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length. This approach presents empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size.
This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale. We propose a two‑parameter model
$E(x)=a\ln\!\bigl(1+bx\bigr),\quad a,b>0,$
that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step. The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function.
It improves interpretability, fairness, and Inter-Rater Reliability (IRR) across both human and AI-generated translations. By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation. The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation congruent with human judgment. Implementation in CAT/LQA systems and implications for human and AI-generated text evaluation are discussed.
@@ 1 Introduction
Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision. It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014], with proper training of annotators. Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment.
@@
In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample. This model typically normalizes errors per 1000 or 2000 words to compare them against a client-defined threshold, e.g., no more than five minor errors per 1000 words. The normalized error density is then used to make a pass-or-fail decision or compute a quality score.
This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words). However, problems arise when evaluating smaller or larger samples.
@@
Clients, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality. Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability. This suggests that perceived translation quality is not a linear function. Evidence from other domains points in the same direction. In speech and video quality, international standards report top-end ’compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic]. Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ’showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450]. Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021]. And in usability engineering, a single ’catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994]. Together, these applied literatures reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
A recent quote from a large institutional client illustrates this point:
’Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them. We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient. The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula. For example, we might feel that if a translation sample is about one page, a single major mistranslation error is already enough to judge it as failing; in a seven-page sample, however, seven such errors are far more than we would be willing to tolerate. Instead, we would prefer to fail the sample already at three or four major errors. This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.’
Another client confirmed these findings and chose to avoid the issue by only evaluating samples of exactly 1000 words. However, in real-world operations, such precisely consistent evaluation sample sizes are rare. Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
@@
Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance. Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations.
@@
This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length. The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases. We capture this pattern with a logarithmic dependency.
Earlier work [GladkoffEtAl2024] introduced a logarithmic scorecard. This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory); we then provide a simple, easily calibrated model for use in current evaluation tools.
@@ 2.1 Analytic TQE and MQM
Analytic Translation Quality Evaluation (TQE) refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications. Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories.
The most prominent and widely adopted framework for such structured annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014]. MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling. Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the client’s preference.
Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict. The MQM framework is currently the most detailed and accurate approach to analytic translation-quality evaluation in professional practice. Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score. For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words.
Refinements to either component therefore directly advance the state of the art in translation-quality measurement.
@@ 2.2 Linear Scoring Models and Their Limits
The dominant approach to scoring in translation quality evaluation is the linear model. Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size (typically 1,000–2,000 words) so that assessments are comparable across jobs of different lengths. For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
Although this proportional scaling appears objective, it fails however to reflect how human content users actually perceive quality. Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450]; however, this tolerance wanes as the document length grows further and the errors begin to compound. For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance. Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
@@
  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a client-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another linear model is needed; a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
@@
When linear scoring is applied indiscriminately across all ranges, it produces systematic friction. Clients and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples. Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment. Because the width of a confidence interval shrinks roughly like $1/\sqrt{n}$—with an additional finite-population correction $\sqrt{(N-n)/(N-1)}$ when the evaluated sample is a large share of the document—variation in the sampled word count can change those intervals considerably. [^2] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
@@ 3 Empirical Evidence of Non-Linearity
To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large enterprise clients. Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
@@ 3.1 Survey Design
Importantly, specific measures were taken to neutralize the influence of the standard linear model, which respondents are accustomed to using in their daily work. Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds. To counteract this bias, the clients’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
@@
  • Intuitively, how many minor errors would you accept in a 1-page sample?
  • Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages?
  • For a 3-page text, how many minor errors seem tolerable—just based on your experience?
  • Don’t calculate—just estimate: how many minor errors would you allow in 4 pages?
  • Based on gut feeling, how many minor errors would still be okay in a 5-page sample?
  • Without applying your formal model, what number of minor errors seems acceptable in 10 pages?
  • Without referencing formulas, what minor error count would intuitively feel acceptable for 20 pages?
@@
While clients used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
@@ 3.2 Results
The most striking finding was a clear non-linear pattern in error tolerance. Freed from the habit of linear extrapolation, respondents indicated that while they would accept more total errors in longer texts, their tolerance grew far more slowly than the sample size. Instead, the tolerance curve rose steeply at first and then gradually flattened. This pattern was consistent across all three clients, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types.
Figures 13 illustrate the raw responses. For each, the horizontal axis represents the sample size in pages, and the vertical axis shows the maximum number of minor errors that the respondent deemed acceptable. In one case, both minor and major error series were collected.
@@
Figure 1: Client 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
@@
Figure 2: Client 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
@@
Figure 3: Client 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
None of the clients’ tolerance curves followed a straight line. Although each client had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve.
The respondents themselves were surprised by this result. When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ’acceptable’ was non-linear. They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate.
This phenomenon aligns with everyday experience. In very short samples, a single error can be jarring and dominate the user’s entire impression of quality. In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness.
@@
This degradation impacts more than just a pleasant "reading experience"; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly. Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
Consistent with Fig. 7 in [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span.
@@
Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size. The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions.
From Figures 1–3 it is already clear that the data points do not fall on a straight line. The logarithmic curve, in contrast, follows the shape of the expert judgments closely.
Quantitative comparison of linear vs. logarithmic fits (technical note)
In other words, the visual impression is also statistically decisive.
Below is the quantitative comparison of linear vs. logarithmic fits on the elicited points.
To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes. This is what the values in Table 1 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better). Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC. This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can. In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
Table 1: Goodness-of-fit statistics for the Client 1 minor-error dataset ($x$ in pages; $E$ in minor points). Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit. Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).

### v3

@@ Abstract
Analytic Translation Quality Evaluation (TQE), based on Multidimensional Quality Metrics (MQM), has traditionally used a linear error-to-penalty scale calibrated to a reference sample size of 1000100020002000 words. However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition. Building on our previously introduced Multi-Range framework, this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length. We present empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size. This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale. We propose a two‑parameter model that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step. The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function. It improves interpretability, fairness, and Inter-Rater Reliability (IRR) across both human and AI-generated translations. By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation. The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation, congruent with human judgment.
keywords:
Translation Quality Evaluation, Statistical Quality Measurement, Translation Evaluation Model, Non-Linear Scoring, Empirical-Based Design
@@ 1 Introduction
Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision. It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014, GladkoffEtAl2024], with proper training of annotators. Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment. In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample. This model typically normalizes errors per 1000 or 2000 words to compare them against a user-defined threshold, e.g., no more than five minor errors per 1000 words. The normalized error density is then used to make a pass-or-fail decision or compute a quality score  [^1] Logrus Global Statistics https://logrusglobal.substack.com/archive . This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words). However, problems arise when evaluating smaller or larger samples [gladkoff2021].
@@
Industrial users, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality. Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability. This suggests that perceived translation quality is not a linear function. Evidence from other domains points in the same direction. In speech and video quality, international standards report top-end ‘compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic]. Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ‘showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450]. Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021]. In usability engineering, a single ‘catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994]. Together, these applied literature reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
@@
Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance. Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations. [^2] We list some real-world institutional users’ experiences on translation evaluation in Appendix A.
@@
This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in our earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length. The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases. We capture this pattern with a logarithmic dependency. The Multi-Range TQE work introduced a logarithmic scorecard. This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory). We also provide a simple, easily calibrated model for use in current evaluation tools. The structure of this paper is organized as below: Section 2 introduces the background on analytical TQE and MQM, Section 3 presents our empirical evidence of non-linearity, Section 4 offers psychophysical and cognitive foundation, Section 5 details the non-linear scoring model, Section 6 and 7 cover the practical implications and step-by-step guide on building a non-linear scorecard, followed by Limitations and Future Work (Section 8).
@@ 2.1 Analytic TQE and MQM
Analytic TQE refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications. Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories. The most prominent and widely adopted framework for such structured annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014]. MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling. Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the user’s preference.
Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict. Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score. For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words  [^3] https://themqm.org/.
@@ 2.2 Linear Scoring Models and Their Limits
The dominant approach to scoring in translation quality evaluation is the linear model. Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size so that assessments are comparable across jobs of different lengths. For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
Although this proportional scaling appears objective, it fails to reflect how human content users actually perceive quality. Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450]. However, this tolerance wanes as the document length grows further and the errors begin to compound. For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance. Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
Building on our earlier work on Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
@@
  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a user-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another model is needed: a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
@@
When linear scoring is applied indiscriminately across all ranges, it produces systematic friction. Users and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples. Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment. Because the width of a confidence interval shrinks roughly like $1/\sqrt{n}$—with an additional finite-population correction $\sqrt{(N-n)/(N-1)}$ when the evaluated sample is a large share of the document—variation in the sampled word count can change those intervals considerably. [^5] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
@@ 3 Empirical Evidence of Non-Linearity
To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large institutional users (anonymized). Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
@@ 3.1 Survey Design
Importantly, specific measures were taken to neutralize the influence of the standard linear model, which respondents are accustomed to using in their daily work. Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds. To counteract this bias, the institutions’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
@@
Intuitive Error Tolerance Questions Intuitively, how many minor errors would you accept in a 1-page sample? Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages? For a 3-page text, how many minor errors seem tolerable—just based on your experience? Do not calculate—just estimate: how many minor errors would you allow in 4 pages? Based on gut feeling, how many minor errors would still be okay in a 5-page sample? Without applying your formal model, what number of minor errors seems acceptable in 10 pages? Without referencing formulas, what minor error count would intuitively feel acceptable for 20 pages?
@@
While institutions used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
@@ 3.2 Results
The most striking finding was a clear non-linear pattern in error tolerance. Freed from the habit of linear extrapolation, respondents indicated that while they would accept more total errors in longer texts, their tolerance grew far more slowly than the sample size. Instead, the tolerance curve rose steeply at first and then gradually flattened. This pattern was consistent across all three anonymized institutions, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types. Figures 13 illustrate the raw responses. For each, the horizontal axis represents the sample size in pages, and the vertical axis shows the maximum number of minor errors that the respondent deemed acceptable. In one case, both minor and major error series were collected.
@@
Figure 1: Institution 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
@@
Figure 2: Institution 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
@@
Figure 3: Institution 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
@@
None of the institutions’ tolerance curves followed a straight line. Although each organization had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve. The respondents themselves were surprised by this result. When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ‘acceptable’ was non-linear. They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate. This phenomenon aligns with everyday experience. In very short samples, a single error can be jarring and dominate the user’s entire impression of quality. In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness. This degradation impacts more than just a pleasant ”reading experience”; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly. Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
@@
Consistent with Fig. 7 in our pilot study from [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span. Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size. The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions. From Figures 13, it is already clear that the data points do not fall on a straight line. The logarithmic curve, in contrast, follows the shape of the expert judgments closely. For readers who want the formal evidence behind Fig. 3: we detail the linear vs logarithmic fits in Appendix B.